# Scrapers selectively respect robots.txt directives: evidence from a large-scale empirical study

### v1

@@ 4.3. Results and Discussion
We begin with RQ1. In Table 5, we bold the directive with the highest compliance rate for each bot. From this, we see that $5$ categories of bots comply the most with crawl delay, while $2$ categories of bots each have highest compliance with endpoint access and disallow all. Furthermore, we compute the average compliance rate for each directive, the last row in the table, and find the highest average compliance with crawl-delay—nearly 2x that of the next-highest, endpoint access. Based on this analysis, we make the following conclusion:
@@
Figure 9. Many bots comply by default with the 30 second crawl delay (leftmost graph), but the rate of compliance decreases overall as robots.txt restrictions tighten. We exclude results for exempted SEO bots and highlight statistically significant ($p\leq 0.05$) shifts with red dotted lines.
@@
Apache-httpclient Apache Other Unknown 0.130 0.043 0.000
@@
Bytespider ByteDance AI Data Scraper No 0.398 0.000 0.028
@@
Facebookexternalhit Meta Fetcher No 0.923 0.276 0.375
@@
PerplexityBot Perplexity AI Search No 0.933 0.900 0.202
@@
Python-requests Open Source Other Unknown 0.462 0.051 0.007
@@
SkypeUriPreview Microsoft Other Yes 0.726 0.010 0.000
@@
Methodology. To understand the frequency with which bots check robots.txt , we first find additional sites in our dataset for which robots.txt were deployed with meaningful restrictions and were publicly available. Three out of the $36$ total sites had robots.txt that met this criteria. All three files were identical and implemented simple restrictions on /404 and /secure endpoints on their respective sites. Through dataset inspection, we found that all bots respected these restrictions.
@@
We then grouped these results by bot category (e.g. Dark Visitors categories used previously), to give a better sense of aggregate bot behavior, and report the proportion of bots in each category that check robots.txt within each of the $5$ given time windows in Figure 10.
Results. Some categories of bots (scrapers, archivers, and intelligence gatherers) consistently re-check robots.txt within $12$ hours if they check it at all. Other categories exhibit more variable behavior, but increasing the window of time between robots.txt checks typically leads to a higher re-check rate. AI-related bots (AI Assistants and AI search crawlers) have the lowest re-check rates with less than $40\%$ of bots checking robots.txt within a $168$ hour window. Thus, frequent updates to robots.txt to prevent unwanted scraping may not effectively prevent scraping, particularly for AI bots.
@@ 5.2. User agent spoofing
Motivation. While running our analysis in §4, we observed that a few bots had high-but-not perfect compliance rates. For example, Googlebot had a $0.65$ compliance rate with the 30 second crawl delay in robots.txt , despite checking robots.txt several times during this experiment. Similarly, Amazonbot had a $0.96$ compliance rate, even though it too regularly checked robots.txt . Several other high profile bots like GPT-Bot and bingbot exhibited similar trends.
This led us consider whether bot “spoofing” played a role in lowering measure compliance rate for certain well-known bots. The useragent field in a web access request is manually set by the requesting agent, so malicious bots can easily manipulate this field to misrepresent themselves, potentially circumventing certain robots.txt restrictions. Some bots, like Googlebot, have elevated privileges in many robots.txt files (including our experiments). By spoofing Googlebot’s user agent, a bot could fly under the radar for longer before being banned. They could eventually be detected via IP address checks and behavioral patterns, but spoofing could help them avoid early flagging.
@@
Methods. We analyzed web traffic data for known bots and observed empirically that most bots are overwhelmingly associated with a particular ASN. For example, $99.62\%$ of Googlebot’s traffic comes from the GOOGLE-CLOUD-PLATFORM ASN, while $99.76\%$ of GPT-Bot’s traffic is from MICROSOFT-CORP. Based on this, we develop an empirical heuristic that if a bot’s traffic is associated $90\%$ of the time with one ASN, other ASNs associated with this user agent are likely spoofed. Using this heuristic, we analyze web accesses from our dataset and flag any bots associated with more than $1$ ASN for which one ASN has over $90\%$ of the traffic. We focus specifically on well-known bots, for which spoofing is more interesting.
Bot Main ASN ($>90\%$ of accesses) Possible spoofing ASNs ($<5\%$ of accesses)
@@
Results. Overall, we identify $18$ bots for which spoofing may have occurred. In Table 8, we show the flagged bots, the dominant ASN (making up $\geq 90\%$ of bot traffic), and the suspicious ASNs ($\leq 1\%$ of traffic). Although many flagged bots are only associated with two ASNs, some popular bots such as Googlebot are associated with up to $24$. Our analysis cannot prove definitively that these infrequent ASNs belong to spoofers (maybe Google has a contract with Telefonica_de_Espana?), but it strongly suggests this is the case.
@@ 6. Discussion
From this large-scale study, we note several limitations of using robots.txt as a deterrent for web scrapers, particularly AI scrapers. First, we observe that bots are less likely to comply with more restrictive robots.txt directives, such as denying access to certain subpages, but are friendlier about respecting crawl delay. This implies that leveraging robots.txt alone to prevent unwanted scraping could be ineffective. Second, we observed that many bots with known user agents in our data never checked the robots.txt file or check it infrequently. This means that any changes made to robots.txt to prevent unwanted scraping (e.g. (Longpre et al., 2024)) would not be noticed by the scraper for significant time (if they are respected at all). Finally, we observe numerous instances of potential spoofing of well-known bots, indicating that malicious user-agents may attempt to evade robots.txt restrictions by masquerading as bots with higher privileges. Together, these suggest that robots.txt does not provide fine-grained access control over web content, highlighting the need for more strongly-enforceable methods to prevent unwanted scraping.
@@
Appendix A Statistical tests for change in bot compliance
@@
ChatGPT-User -2.07 3.86e-02 -0.18 8.60e-01 20.62 0.00e+00
Claudebot 0.44 6.62e-01 8.95 0.00e+00 8.24 2.22e-16
@@
Go-http-client 14.05 0.00e+00 19.24 0.00e+00 5.23 1.72e-07
GPTBot 11.54 0.00e+00 7.43 1.09e-13 24.20 0.00e+00
@@
Python-requests 11.46 0.00e+00 5.04 4.64e-07 1.83 6.71e-02
SemanticScholarBot 10.39 0.00e+00 6.86 6.99e-12 7.36 1.81e-13
SemrushBot 0.91 3.62e-01 9.90 0.00e+00 10.18 0.00e+00

### v2

@@ 4.3. Results and Discussion
We begin with RQ1. In Table 5, we bold the directive with the highest compliance rate for each bot. From this, we see that $5$ categories of bots comply the most with crawl delay, while $2$ categories of bots each have highest compliance with endpoint access and disallow all. Furthermore, we compute the average compliance rate for each directive, the last row in the table, and find the highest average compliance with crawl-delay—nearly 2x that of the next-highest, endpoint access. Based on this analysis, we conclude that:
@@
Figure 9. Many bots comply by default with the 30 second crawl delay (leftmost graph), but the rate of compliance decreases overall as robots.txt restrictions tighten. We exclude results for exempted SEO bots and highlight statistically significant ($p\leq 0.05$) shifts with red dotted lines. We also exclude results for bots that appear to have spoofed their user agent. Results for only putative spoofed bots are in Appendix A.2, and additional analysis of observed spoofing behavior is in §5.
@@
Apache-httpclient Apache Other Unknown 0.091 0.043 0.000
@@
Bytespider ByteDance AI Data Scraper No 0.398 0.000 0.000
@@
Facebookexternalhit Meta Fetcher No 0.920 0.281 0.375
@@
PerplexityBot Perplexity AI Search No 0.933 0.897 0.202
@@
Python-requests Open Source Other Unknown 0.462 0.051 0.000
@@
SkypeUriPreview Microsoft Other Yes 0.726 0.000 0.000
@@
Methodology. To understand the frequency with which bots check robots.txt , we first find additional sites in our dataset for which robots.txt were deployed with meaningful restrictions (e.g. contained at least one directive designed to change bot behavior) and were publicly available. Three out of the $36$ total sites had robots.txt that met this criteria. All three files were identical and implemented simple restrictions on /404 and /secure endpoints on their respective sites. Through dataset inspection, we found that all bots respected these restrictions.
@@
We then grouped these results by bot category (e.g. Dark Visitors categories used previously), to give a better sense of aggregate bot behavior, and report the proportion of bots in each category that check robots.txt within each of the $5$ time windows in Figure 10.
Results. Some categories of bots (scrapers, archivers, and intelligence gatherers) consistently re-check robots.txt within $12$ hours if they check it at all. Other categories exhibit more variable behavior, but increasing the window of time between robots.txt checks typically leads to a higher re-check rate. AI-related bots (AI Assistants and AI search crawlers) have the lowest re-check rates with less than $40\%$ of bots checking robots.txt within a $168$ hour window. Thus, frequent updates to robots.txt may not effectively prevent scraping, particularly for AI bots.
@@ 5.2. User agent spoofing
Motivation. Our analysis of §4 included checks for user agent “spoofing.” Here, we describe additional details of our methods for detecting spoofing and results. A key motivation for this investigation was our observation that a few bots had high-but-not perfect compliance rates. For example, Googlebot had a $0.65$ compliance rate with the 30 second crawl delay in robots.txt , despite checking robots.txt several times during this experiment. Similarly, Amazonbot had a $0.96$ compliance rate, even though it regularly checked robots.txt . Several other popular bots like GPT-Bot and bingbot had similar trends.
This led us consider how much bot “spoofing” affected measured compliance rate for certain well-known bots. The useragent field in a web access request is manually set by the requesting agent, so malicious bots can easily manipulate this field to misrepresent themselves, potentially circumventing certain robots.txt restrictions. Some bots, like Googlebot, have elevated privileges in many robots.txt files (including our experiments). By spoofing Googlebot’s user agent, a bot could fly under the radar for longer before being banned. They could eventually be detected via IP address checks and behavioral patterns, but spoofing could help them avoid early flagging.
@@
Methods. We analyzed web traffic data for known bots and observed empirically that most bots are overwhelmingly associated with a particular ASN. For example, $99.62\%$ of Googlebot’s traffic comes from the GOOGLE-CLOUD-PLATFORM ASN, while $99.76\%$ of GPT-Bot’s traffic is from MICROSOFT-CORP. Based on this, we develop an empirical heuristic that if a bot’s traffic is associated $90\%$ of the time with one ASN, other ASNs associated with this user agent are likely spoofed. Using this heuristic, we analyze web accesses from our dataset and flag any bots associated with more than $1$ ASN for which one ASN has over $90\%$ of the traffic. Here, we focus specifically on well-known bots, for which spoofing is more interesting.
Bot Main ASN ($90\%$ of accesses) Possible spoofing ASNs ($5\%$ of accesses)
@@
Results. Overall, we identify $18$ bots for which spoofing may have occurred. In Table 8, we show the flagged bots, the dominant ASN (making up $\geq 90\%$ of bot traffic), and the suspicious ASNs ($\leq 1\%$ of traffic). Although many flagged bots are only associated with two ASNs, some popular bots such as Googlebot are associated with up to $24$. Our analysis cannot prove that these infrequent ASNs belong to spoofers (maybe Google contracts with Telefonica_de_Espana?), but it strongly suggests this is the case.
@@
Limitations. The study of spoofing presented here has several limitations. First, our threshold of $90\%$ of traffic originating from a single ASN is somewhat arbitrary. Future work could explore whether adjusting this threshold surfaces other interesting behaviors. Second, our analysis does not allow us to definitively state whether a bot is spoofing its user agent. Future work could consider alternative means, such as honeypots, that could enable more confident conclusions about which bot instances are spoofed.
@@ 6. Discussion
From this large-scale study, we note several limitations of using robots.txt as a deterrent for web scrapers, particularly AI scrapers. First, we observe that bots are less likely to comply with more restrictive robots.txt directives, such as denying access to certain subpages, but are friendlier about respecting crawl delay. This implies that leveraging robots.txt alone to prevent unwanted scraping could be ineffective. Second, we observed that many bots with known user agents in our data never checked the robots.txt file or check it infrequently. This means that any changes made to robots.txt to prevent unwanted scraping (e.g. (Longpre et al., 2025)) would not be noticed by the scraper for significant time (if they are respected at all). Finally, we observe numerous instances of potential spoofing of well-known bots, indicating that malicious user-agents may attempt to evade robots.txt restrictions by masquerading as bots with higher privileges. Together, these suggest that robots.txt does not provide a universally respected signal for bot behavior, highlighting the need for more strongly-enforceable methods to prevent unwanted scraping.
Limitations of our study. This study has several drawbacks that may limit the generalizability of its findings. The first study only considered traffic from $36$ websites owned by our institution. While these websites saw significant traffic during the course of data collection, they represent a small subset of sites online, and bots may behave differently on different types of sites (e.g. sites not affiliated with a large, well-resourced university). The second robots.txt study was conducted on a single website, raising the possibility that bots behaved differently on this site than elsewhere. Both studies only capture a snapshot of bot behavior during a 45 day time window. Bot behavior will likely evolve over time, making it difficult to draw permanent lessons from our findings. Finally, in the first study, bot behavior was not static over time, which could impact findings from the second study.
@@
Appendix A Appendix
A.1. Ethics
This paper raises few ethical concerns, besides those associated with potentially private data included in our user study. As specified in the paper, data collection and analysis adhered to a study protocol approved by our local Institutional Review Board (IRB). We followed all IRB guidelines for anonymizing and storing this dataset, including making significant efforts to exclude non-bot behavior from the dataset and preserve user privacy. We removed sessions originating from logged-in users from our institution or from institution-hosted WiFi networks, and anonymized IP addresses by hashing them. All data was stored and analyzed on secure servers only accessible by the research team.
A.2. Spoofing analysis from §4.3
Our §4.3 analysis of bot compliance with robots.txt  excluded bots with apparently spoofed user agents. We used the methodology presented in §5 to identify potential spoofing bots: we look at bots for which $90\%$ of traffic originates from a single ASN, and flag as possibly spoofed instances any requests with the same useragent originating from another ASN. In Figure 11, we present parallel results to Figure 9 but for putative spoofed bots.
Directive Crawl delay Endpoint access Disallow all
Legitimate requests 10333 11397 8859
Potentially spoofed requests 60 36 247
Table 9. Potentially spoofed requests account for less than $0.1\%$ of all traffic under our three crawl delay directives. We report the number of perceived legitimate and potentially spoofed requests recorded during each of the three robots.txt experiments.
Overall, there were few instances of potentially spoofed requests flagged by our heuristic—see Table 9. However, for the spoofed bots we did observe, there was little respect for robots.txt  directives. Particularly for the endpoint access and disallow all directives, spoofed bot behaviors typically did not change before/after the directive was implemented. The two notable exceptions to this were PerplexityBot under the endpoint access directive and Bytespider under the disallow all directive. Both bots significantly changed their behavior, echoing the behavior seen in presumably true instances of the bots in Figure 9. This perhaps indicates that these were not spoofed instances, but rather instances of the true bot leveraging a different ASN than typically seen. Future work is needed to be able to definitively identify bots with spoofed user agents and avoid such ambiguity.
Figure 11. The majority of potentially spoofed bots respond less—if at all—to robots.txt  changes than their non-spoofed counterparts (see Figure 9). The two exceptions are PerplexityBot and Bytedance bots under the Endpoint Access and Disallow All directives, which may be misidentified by our spoofing detection heuristic. We highlight statistically significant ($p\leq 0.05$) shifts with red dotted lines.
A.3. Statistical tests for change in bot compliance
@@
ChatGPT-User -2.07 3.86e-02 -0.18 8.60e-01 20.62 0.00e+$0$
Claudebot 0.44 6.62e-01 8.95 0.00e+$0$ 8.24 2.22e-16
@@
Go-http-client 14.05 0.00e+$0$ 19.24 0.00e+$0$ 5.23 1.72e-07
GPTBot 11.54 0.00e+$0$ 7.43 1.09e-13 24.20 0.00e+$0$
@@
Python-requests 11.46 0.00e+$0$ 5.04 4.64e-07 1.83 6.71e-02
SemanticScholarBot 10.39 0.00e+$0$ 6.86 6.99e-12 7.36 1.81e-13
SemrushBot 0.91 3.62e-01 9.90 0.00e+$0$ 10.18 0.00e+$0$